OpenStudy (anonymous):

solve for x and show steps: log(x-2)-logx=3

OpenStudy (amistre64):

we solved for x in the other question; its 8 .... right?

OpenStudy (amistre64):

log is 10 base here right?

OpenStudy (anonymous):

yeah log base 10

OpenStudy (amistre64):

\[10^{(log(x-2)-log(x))}=10^3\] \[\frac{10^{log(x-2)}}{10^{log(x)}}=10^3\] \[\frac{x-2}{x}=1000\]

OpenStudy (amistre64):

\[\frac{x-2}{x}=1000\] \[\frac{x}{x}-\frac{2}{x}=1000\] \[1-\frac{2}{x}=1000\] \[-\frac{2}{x}=999\] i think its right, but that looks peculiar

OpenStudy (anonymous):

thats what i got but i wasnt sure. i did it a different way too

OpenStudy (amistre64):

its about -.104791 if it is

OpenStudy (amistre64):

different ways are good, i tend to try to show that math aint got to be done in one specific move

OpenStudy (anonymous):

when you plug x back in, does it make the equation equal?

OpenStudy (amistre64):

cant tell yet :) i havent tried to compute it

OpenStudy (amistre64):

log(x-2)/x =3 10^log(x-2)/x =10^3 (x-2)/x = 1000 x-2 = 1000x -2 = 999x x = -2/999

OpenStudy (amistre64):

ugh, cant have x = a negative number for starters

OpenStudy (anonymous):

true i forgot about that

OpenStudy (amistre64):

but wolfram gives the answer as a negative too , so i dunno

OpenStudy (anonymous):

would it be no solution?

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=log%28x-2%29-log%28x%29+%3D+3&a=*FunClash.log-_*Log10.Log- no were right, somehow; but wolf was reading it as ln, had to tell it to do log

OpenStudy (anonymous):

okay so is the answer x=-2/999 or no solution?

OpenStudy (amistre64):

im going with -2/999

OpenStudy (anonymous):

alright thanks